OpenStudy (anonymous):

What is the approximate circumference of a circle with a diameter of 34 m? (Use π = 3.14.) A. 106.76 m B. 213.52 m C. 427.04 m D. 53.38 m

OpenStudy (anonymous):

@timaashorty

OpenStudy (anonymous):

do you need the formula for circumference?

OpenStudy (richyw):

\[\pi\equiv \frac{\text{circumference}}{\text{diameter}}\]

OpenStudy (richyw):

so in this case the exact circumference is \(34\pi\)

OpenStudy (richyw):

(in meters)

OpenStudy (timaashorty):

Sorry I was off keyboard this whole time! So, What @richyw and @jeremyggg said are correct. @OwlGirl212 use \[34\pi\] and you will get your circumference. \[34 \times 3.14 = 106.76 m \] So your answer is \[\huge A. \]

OpenStudy (anonymous):

ok thx for for late reply